OpenStudy (anonymous):

Express cos^2 x in terms of cos2x

OpenStudy (anonymous):

cos2x = 2 cos^2 x - 1

OpenStudy (ghazi):

\[\cos^ x=\frac{ (1+ \cos 2x) }{ 2 }\]

OpenStudy (ghazi):

* cos^2 x

OpenStudy (ghazi):

in the left side

OpenStudy (anonymous):

Wow that was really quick. Thanks to both of you, Im going to look it over now.

OpenStudy (ghazi):

:) YW

OpenStudy (anonymous):

Thanks, I understand it now.

OpenStudy (ghazi):

:)